Subject: APO 3.5-m Users Committee minutes 5/10/04

From: Bruce Gillespie

Submitted: Thu, 13 May 2004 12:18:33 -0600

Message number: 778 (previous: 775, next: 779 up: Index)

					APO 3.5-m Users Committee Phonecon, 5/10/04

Attending:  Rene Walterbos, Michael Strauss, Jon Holtzman, Ed Turner, 
Bruce Gillespie, Russet McMillan, Karl Glazebrook

Absent:  Al Harper, Don York, Bruce Balick, Jim Green, John Bally

Minutes taken by Bruce Gillespie


NIC-FPS report:

Nobody on the telecon was able to give an update on NIC-FPS status, so 
Ed Turner later asked the team to provide a written report, which 

Hi Ed,
Here's a brief update on NIC-FPS.

Detector--getting images on the Leach controller and are continuing to 
optimize.  The Rockwell controller work is on hold for now.

Filter wheels--functional tests proceeding well, both warm and cold.  
Repeatability tests are producing good results.

Etalon mover--initial tests of mechanism (warm and cold) were 
acceptable, but additional work is in progress to improve results. 
Machining of the etalon mounting hardware is in progress.

Etalon and controller--CS-100 testing at vendor is complete and 
controller is being shipped from U.K.

Dewar--Aero-glazing is finished and pump-down for a contamination check 
showed no offgasing. MLI blanket construction is planned for this week.

Optical bench--integration is underway, including mounting all 
components, making cables, alignment, etc.

Software--preparing control software for integration testing of filter 
wheels and etalon mover.  Much work needs to be done in this area.

Test set-up--all optical components are in hand and assembly/alignment
is underway.

Hope that's sufficient detail to get you up to speed.
Fred [Hearty]


CorMASS tests at APO:

Bruce Gillespie reported that the recent The CorMASS test run at the 
3.5-m telescope last week was quite successful.  Sample spectra were 
taken of many targets that had been suggested by nearly a dozen ARC 
astronomers.  Mike Skrutskie and John Wilson, the Co-PIs for CorMASS, 
were able to take advantage of an easy integration to our systems, plus 
good skies, good seeing, and good pointing to put CorMASS through its 
paces on the telescope.  They report that they were easily able to 
acquire spectra of 15th magnitude objects, and that some examples of 
quickly reduced spectra can be seen at
http:/  These spectra 
represent a small fraction of the science targets for the run.


CorMASS availability in 3Q2004:

Ed Turner reported that Mike Skrutskie and John Wilson have agreed to 
temporarily leave the CorMASS instrument at APO, where it is being kept 
cold and ready to use.  They have
offered its use to ARC astronomers who want to use it through the 
summer, and possibly into the fall.  Although many of our users may 
have submitted their 3rd quarter proposals, substituting CorMASS for 
present and 3rd quarter programs is a possibility that users should 
explore with their scheduler, in order to take advantage of this 
unanticipated opportunity.  Alternatively, or in addition, users may 
submit Director's Discretionary time requests directly to Ed Turner, in 
the standard format.

Operation of CorMASS at APO is currently only in "service" mode, or by 
users being present at the telescope.  Data reduction guides and 
routines exist, and limited operational and data reduction expertise 
has been developed by the site staff.


Last month's minutes were approved without comment.

Next phonecon will be on Monday, June 14, 2004, at 11:30 AM Eastern Time


APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 778 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to